OpenStudy (anonymous):

6x+4=-39 3x+2y=-15 solve for y. i thought it was 18.. but i need clarifacation with these problems.

OpenStudy (anonymous):

I think you meant 6x+4y = -39 right?

OpenStudy (anonymous):

so sorry. 6x+y=-39. just y.

OpenStudy (mathstudent55):

6x + y = -39 3x + 2y = -15 Multiply the second equation by -3 then add both equations together. 6x + y = -39 -6x - 4y = 30

OpenStudy (mathstudent55):

Add the equations together. The x's will add up to zero and will cancel out, and you'll have an equation with only y. then you can easily solve for y.

OpenStudy (anonymous):

thank you.

OpenStudy (mathstudent55):

you're welcome